# Green Edge AI: A Contemporary Survey

### v1

@@ II-C General Design Principles
To effectively reduce energy consumption at EUDs in edge AI tasks, different approaches can be developed. In order to provide a cohesive structure for the literature review in the upcoming sections, we will first establish the general design principles for energy-efficient edge AI. These principles are derived from the energy breakdown analysis of edge AI systems presented in Section II-B.
@@
Energy Efficiency-oriented Optimization: Shifting the design objective of edge AI systems from solely maximizing intelligence performance (e.g., classification accuracy) to maximizing EE is crucial. Specifically, EE of AI systems can be defined as the amount of intelligence obtained with per Joule of energy consumption [61]. Optimizing edge AI systems for EE necessitates a holistic consideration of all three critical energy components discussed in Section II-B. Given that most edge AI applications involve integrated processes of multiple operations, it is essential to address the interplay between sensing, computation, and communication to strike a decent balance between achievable intelligence and energy consumption. Achieving the vision of green and energy-efficient edge AI requires not only significant algorithmic innovations but also fundamental re-architecting of existing mobile networks to enable the all-round cooperation between EUDs and MEC servers.
@@
Adaptation to System Dynamics: Exploiting the dynamics of edge AI systems can enhance EE through adaptive sensing, computation, and communication. For instance, in the case of a moving smart vehicle, the surroundings constantly change, necessitating adjustments in the sampling frequency and resolution of a vehicular LiDAR camera based on the scene’s informativeness. Regarding communication between EUDs and MEC servers, wireless channels exhibit time-varying signal attenuation. Adapting transmission schemes, such as power control, modulation, and coding, to the instantaneous channel conditions can improve the EE of uplink communication [62]. Furthermore, the unique characteristics of edge AI computations, including workload, latency, and accuracy requirements, vary drastically. Therefore, optimizing processing strategies to achieve maximum energy savings at EUDs is essential [10]. Another dimension of system dynamics in edge AI that should be considered is the task state, which can refer to the progress of a learning task or the difficulty of an inference task.
@@
Trading Intelligence for Greenness & Energy Efficiency: As AI algorithms, such as DNN models, approach their performance limits, the energy consumption required to achieve further intelligence improvement increases super-linearly [21, 63]. Therefore, it is crucial to develop application-specific solutions that can provide just-enough intelligence, rather than blindly adopting state-of-the-art approaches. Besides, the fault-tolerance property of DNN models [64, 65] enables to control the trade-off between intelligence and reduced energy consumption. This can be achieved by identifying what to compute, in addition to optimizing where and how to compute, as typically considered for generic computation tasks. In other words, computations that have minimal impact on intelligence performance can be avoided. Such a principle can also be applied to sensing and communication operations in edge AI systems. However, due to the lack of a universal and tractable characterization of the achievable intelligence and the energy consumption of sensing, computation, and communication, the optimization of their trade-off needs to be task-oriented through offline profiling or online learning.
@@ III Energy-Efficient Data Acquisition for Centralized Edge Learning
Although the computations of model training are offloaded in centralized edge learning, sensing the surroundings (i.e., data sampling) and transmitting the sensed data to the MEC server contribute to most of the on-device energy consumption [67, 68]. In this section, we investigate energy-efficient data acquisition approaches for centralized edge learning, which answer the following three questions: 1) How to perform energy-efficient data sampling (Section III-A); 2) How to transmit these data to the MEC server with minimal energy consumption (Section III-B); and 3) How to improve the learning performance when only a small set of training data can be acquired by the MEC server (Section III-C).
@@ III-A Adaptive Sampling
As model training tasks are data-intensive, energy consumed by the sensing modules of EUDs to collect a vast amount of environmental measurements is significant. Wireless sensor networks (WSNs) are typical venues where data are periodically sampled by the distributed sensor nodes. According to the Nyquist Theorem [69], the data sampling rate should be sufficiently high to ensure artifact-free sensing. However, constant-rate sampling fails to exploit the time-varying property of the sensed data and easily results in redundant data samples. In contrast, adaptive sampling, which reacts to behaviors of the sensed data by adjusting the sampling rate in runtime, is effective in reducing the sampling energy cost [68]. For centralized edge learning, it also saves the communication energy since less amount of data needs to be transmitted.
@@ III-B Learning-aware Data Sample Transmission
Albeit some redundant data can be eliminated using adaptive sampling strategies, the remaining ones exhibit different importance to edge learning tasks, bringing ample opportunities for further transmission energy reduction. First, data samples can be prioritized and transmitted in order so the less important ones can be discarded [47, 75]. Second, data samples can be transmitted in reduced quality via lightweight pre-processing such as lossy compression [76]. For both methods, balancing the communication cost and learning performance is crucial.
@@
Data sample importance typically changes at different stages of the learning process. To obtain such information at EUDs, the MEC server needs to feedback the newest learned model, denoted as $\mathbf{w}$, for data importance evaluation [47, 75, 76]. In [76], the importance of a sample $\mathbf{x}$ is determined by the training loss $l\left(\mathbf{x};\mathbf{w}\right)$, of which, a larger value implies higher significance. However, re-calculating the training loss for a large number of data samples whenever the model is updated is both time- and energy-consuming. One solution is to restrict the evaluation on a small subset of training data. It was validated on the MNIST dataset that, only $2.5\%$ of the data samples need to be evaluated to maintain similar learning accuracy. Data importance evaluation is also closely related to the burgeoning area of active learning [77], where prediction uncertainty of a data sample under the newest learned model is commonly adopted [47, 75]. Besides, based on the heuristics that training with highly disparate data samples may produce better models, the Euclidean distance between a data sample of interest and centroid of all available samples at the MEC server, was used as the data importance metric in [78]. The concept of data sample importance was also leveraged for edge robotic systems in [79], where a lightweight HARVESTNET was developed to determine whether to cache and transfer a training sample to a remote server for model re-training.
@@
In contrast to conventional wireless communication systems, energy-efficient transmission schemes for centralized edge learning should consider the data sample importance. In [80], a multi-criteria training data subset (MCTS) algorithm for prioritizing data transmission in a predictive maintenance application was developed, where the high-quality data samples are protected against wireless signal attenuation with more resources. Besides, data compression before transmission was proposed for centralized edge learning, where the compression ratio is determined according to the training loss in  [76] and via an error-bounded compression scheme in [81]. To avoid the complex baseband processing at EUDs, analog wireless transmission was exploited in [47, 75, 78], where data samples are directly mapped to channel symbols and thus vulnerable to noise corruption. As a result, a retransmission mechanism was developed in [75] to eliminate negative effects of noise, where a data sample is retransmitted until the effective signal-to-noise ratio (SNR) is above a threshold. With multiple EUDs, it is also necessary to schedule the most appropriate ones to transmit, which minimizes the communication overhead for a target learning performance [47, 78].
@@
Apart from sample-level importance, the amount of data used for model training also has a great impact on the learning performance. Although the exact relationship between data volume and training loss is hard to find, data-driven methods are able to establish power-law models for approximiate characterizations [82]. By adopting a nonlinear function mapping the data volume to the training loss, a learning-centric power allocation problem for centralized edge learning was solved in [83] to decide the amount of data samples used for the training of multiple classification models. This study was extended in [84] and [85], respectively incorporating wireless resource allocation and multi-device scheduling. More realistic scenarios accounting for continuous data arrivals and limitations of MEC servers were investigated in [86, 87].
@@ III-C Centralized Edge Learning with Small Datasets
1) Data Augmentation: Data augmentation (DA) refers to the process of enhancing the size, quality, and diversity of a training dataset [88], hoping to obtain a more accurate model than using the original dataset. For instance, basic spatial operations such as rotation and noising, can be applied to images for DA. In addition to hand-crafted methods, generative adversarial network (GAN) based image synthesis [89] becomes a new favourite to increase the robustness and generalizability of DA [90, 91]. DA methods for other data types are less investigated, despite with recent advances on speech recognition [92], video analytics [93], and NLP [94].
@@
DA has found its adoptions in a variety of edge AI applications. In [95], the feasibility of generating sequential IoT data via GAN for classification is validated, even with a small volume of training data. The hybrid smart medical architecture for electrocardiogram diagnostics developed in [96] adopts a similar approach to overcome the class imbalance issue in the training dataset. For WiFi signal-based human activity recognition, non-learning based data synthesis was utilized in [97] to mitigate the negative impacts caused by human movement and subject-specific bias. We are aware that while [95, 96, 97] did not directly address the energy consumption of EUDs, DA improves the learned model performance in the presence of a small and possibly low-quality training dataset. In other words, the amount of data transmitted from EUDs to the MEC server and their quality can be reduced to achieve the target learning performance with DA, thereby saving the sampling and communication energy.
2) Transfer Learning: The objective of transfer learning (TL) is to reuse the learned knowledge for a new but related model training task [98]. An attractive benefit of knowledge transfer techniques for centralized edge learning is the reduced requirement of labeled training data [99], therefore minimizing the cost of training data acquisition, including the sensing and communication energy. Besides, as MEC environments and data distributions may experience regular drifts, AI models need be periodically finetuned [100]. This can also be assisted with TL to avoid re-training models entirely from scratch.
The developments of industrial IoT applications have been actively utilizing TL techniques. The use of TL for industrial IoT component recognition was investigated in [101], where a pre-trained VGG-16 network on ImageNet was transferred to a new scenario with a small target domain T-Less dataset of texture-less objects. Compared with training a convolutional neural network (CNN) model from scratch, knowledge transfer reduces both the required training data volume and training time, while attaining better training accuracy. For edge camera image recognition in unmanned stores, the experimental results in [102] showed that with an elite-instance-based matching approach that appropriately selects the images to be used for TL, up to $70\%$ of source image samples can be eliminated from transmission to the target cameras. TL was also utilized for device-free crowd counting based on the WiFi channel state information (CSI) [103]. In situations where data transmission is extremely costly and energy-expensive, some data categories may only have a few samples for model training at the MEC server, for which, few-shot learning (FSL) approaches [104] can be leveraged. Sample applications include object detection [105], sentiment analysis on text [106], and IoT traffic classification [107].
@@ III-D Takeaways
This section summarizes energy-efficient data acquisition methods for centralized edge learning, which were developed following either of the two approaches: 1) To improve the EE of data sampling and transmission; 2) To improve the learning performance with only a small training dataset available at the MEC server. For the former approach, tracable models relating various decision variables, e.g., data sampling frequency, resolution, and context, with the learning performance are essential, which can be obtained via offline data-fitting or online learning. Although their effectiveness has been observed, existing works are mainly based on heuristics without a strong theoretical support. Besides, accurate energy models of sensing modules on EUDs are necessary to formulate valid optimization problems. For cameras, the energy model developed for visual computing systems in [68] can be used. Also, a power model suitable for three-dimensional depth sensors (e.g. Lidar cameras) can be found in [108]. The second approach stems from the field of few-shot learning [104], while other techniques, such as meta-learning-based methods, may also be applied. To further reduce the communication energy consumption, it is important to understand the minimum amount of training data that needs to be transmitted from EUDs to the MEC server with the joint aid of DA and TL techniques. Since the three classes of techniques reviewed respectively in Section III-A, III-B, and III-C were examined separately, their composite effects on achieving energy-efficient data acquisition for centralized edge learning deserve further investigtions.
@@ IV Energy-Efficient Cooperative Training at the Edge Network
Centralizing the model training process at an MEC server necessitates substantial bandwidth to transfer large volumes of training data, a requirement that is not always feasible in wireless edge networks. This approach may also infringe upon data privacy regulations [109], as local data may be private and confidential. However, advancements in embedded hardware technologies have empowered a growing population of EUDs to perform on-device model training [110, 111]. As such, cooperative model training at the edge network, which distributes a model training task among multiple EUDs, has been extensively explored to leverage both the ubiquity of big data and computational resources. In this section, we focus on federated learning (FL) [112], a de facto standard for privacy-preserving cooperative model training. However, the principles of energy-efficient FL design in mobile edge networks are applicable to other cooperative training frameworks [113]. The energy-efficient design approaches to be discussed in this section for cooperative training at the edge network are summarized in Fig. 3.
@@ IV-A Background of Federated Learning
FL was first introduced by Google in 2016 [114]. In this framework, a set of client nodes (e.g., EUDs) collaborate to train an ML model under the coordination of a parameter server (PS), as illustrated in Fig. 4. During each training iteration, clients initially download the global model from the PS. Subsequently, each client independently updates the downloaded model using its local data, typically through methods such as stochastic gradient descent (SGD). Following this, the local model updates are uploaded to the PS for global aggregation. As the training data remains local to the clients, data privacy is significantly better protected compared to the centralized learning paradigm discussed in Section III.
While initial FL systems utilized cloud-based PSs, there is a growing interest in deploying FL at the mobile edge network, a concept known as federated edge learning (FEEL). This approach holds potential for facilitating rapid cooperative training [115], with an MEC server designated as the PS. Given that FEEL operates over resource-constrained mobile edge networks, both the model transmission and local training processes need to be carefully optimized to maximize EE without causing significant degradation to learning performance. In the following sections, we introduce two categories of techniques as potential pathways to realize energy-efficient FEEL, including model and gradient compression (Section IV-B), and resource management (Section IV-C). We discuss these techniques from an EE perspective, although some of them are also essential for communication-efficient FEEL [34].
Figure 4: A typical federated learning system and its training procedures.

### v2

@@ II-C General Design Principles
To effectively reduce energy consumption at EUDs for edge AI tasks, a wide range of approaches can be developed. In order to provide a cohesive structure for the literature review in the upcoming sections, we will first establish the general design principles for energy-efficient edge AI. These principles are derived from the energy breakdown analysis of edge AI systems presented in Section II-B.
@@
Energy Efficiency-oriented Optimization: Shifting the design objective of edge AI systems from solely maximizing intelligence performance (e.g., classification accuracy) to maximizing EE is crucial. Specifically, EE of AI systems can be defined as the amount of intelligence obtained with per Joule of energy consumption [66]. Optimizing edge AI systems for EE necessitates a holistic consideration on all the three critical energy components discussed in Section II-B. Given that most edge AI applications involve integrated processes of multiple operations, it is essential to investigate the intricate interplay among sensing, computation, and communication in order to strike a decent balance between achievable intelligence and energy consumption, which vary drastically for three edge AI tasks. Therefore, achieving the vision of green and energy-efficient edge AI requires not only significant algorithmic innovations with EE-oriented design objectives, but also fundamental re-architecting of existing mobile networks to enable the all-round cooperation between EUDs and MEC servers to best utilize their computational resources.
@@
Adaptation to System Dynamics: Exploiting the dynamics of edge AI systems can enhance EE through adaptive sensing, computation, and communication. For instance, in the case of a moving smart vehicle, the surroundings constantly change, necessitating timely adjustments in the sampling frequency and resolution of a vehicular LiDAR camera based on the scene’s informativeness. Regarding information exchange between EUDs and MEC servers, since wireless channels exhibit time-varying signal attenuation, adapting transmission schemes, such as power control, modulation, and coding, to the instantaneous channel conditions can improve the EE of uplink communication [67]. Furthermore, the unique characteristics of edge AI computations, including workload, latency, and accuracy requirements, vary significantly. Therefore, optimizing processing strategies to achieve maximum energy savings at EUDs is also essential [10]. Another dimension of system dynamics in edge AI that should be considered is the task state, which can refer to the progress of a learning task or the sample difficulty of an inference task. In addition, it is important to monitor the arrivals/departures of EUDs and their requested edge AI services to facilitate energy-efficient dynamic resource allocation.
@@
Trading Intelligence for Greenness & Energy Efficiency: As AI algorithms, such as DNN models, approach their performance limits, the energy consumption required to achieve further intelligence improvement increases super-linearly [21, 68]. Therefore, it is crucial to develop application-specific solutions that can provide just-enough intelligence, rather than blindly adopting state-of-the-art approaches. Besides, the fault-tolerance property of DNN models [69, 70] enables to control the trade-off between intelligence and reduced energy consumption. This can be achieved by identifying what to compute, in addition to optimizing where and how to compute, as typically considered for generic computation tasks. In other words, computations that have minimal impact on intelligence performance should be avoided. Such a principle can also be applied to the sensing and communication operations in edge AI systems. For instance, the data scaling law for model training [71, 72] indicates that increasing the training dataset size beyond a limit has saturating benefit on accuracy. Thus, sensing and communication energy consumption in centralized edge model training can be reduced by collecting just enough amount of data samples for a target model accuracy. However, due to the lack of a universal and tractable characterization of the achievable intelligence and the energy consumption of sensing, computation, and communication, the optimization of their trade-off needs to be task-oriented with prior information obtained through offline profiling or online learning.
@@ III Energy-Efficient Data Acquisition for Centralized Edge Learning
Although the computations of model training are offloaded in centralized edge learning, sensing the surroundings (i.e., data sampling) and transmitting the sensed data to the MEC server contribute to most of the on-device energy consumption [74, 75], which increases with the required amount of training data samples. In this section, we investigate energy-efficient data acquisition approaches for centralized edge learning, which answer the following three questions: 1) How to perform energy-efficient data sampling (Section III-A); 2) How to transmit these data to the MEC server with minimal energy consumption (Section III-B); and 3) How to improve the learning performance when only a small training dataset can be acquired by the MEC server (Section III-C).
@@ III-A Adaptive Sampling
As model training tasks are data-intensive, energy consumed by the sensing modules of EUDs to collect a vast amount of environmental measurements is significant. Wireless sensor networks (WSNs) are typical venues where data are periodically sampled by the distributed sensor nodes. According to the Nyquist Theorem [76], the data sampling rate should be sufficiently high to ensure artifact-free sensing. However, constant-rate sampling fails to exploit the time-varying property of the sensed data and easily results in redundant data samples. In contrast, adaptive sampling, which reacts to behaviors of the sensed data by adjusting the sampling rate in runtime, is effective in reducing the sampling energy cost [75, 77]. For centralized edge learning, it also saves the communication energy since less amount of data needs to be transmitted.
@@ III-B Learning-aware Data Sample Transmission
Albeit some redundant data can be eliminated using adaptive sampling strategies, the remaining ones exhibit different importance to centralized edge learning tasks, bringing ample opportunities for further transmission energy reduction. First, data samples can be prioritized and transmitted so the less important ones can be discarded [52, 83]. Second, data samples can be transmitted in reduced quality via lightweight pre-processing such as lossy compression [84]. For both methods, balancing the communication cost and learning performance is crucial.
@@
Data sample importance typically changes at different stages of the learning process. To obtain such information at EUDs, the MEC server needs to feedback the newest learned model, denoted as $\mathbf{w}$, for data importance evaluation [52, 83, 84]. In [84], the importance of a sample $\mathbf{x}$ is determined by the training loss $l\left(\mathbf{x};\mathbf{w}\right)$, of which, a larger value implies higher significance. However, re-calculating the training loss for a large number of data samples whenever the model is updated is both time- and energy-consuming. One solution is to restrict the importance evaluation on a small subset of training data. It was validated on the MNIST dataset that, only $2.5\%$ of the data samples need to be evaluated to maintain similar learning accuracy. Data importance evaluation is also closely related to the burgeoning area of active learning [85], where prediction uncertainty of a data sample under the newest learned model is commonly adopted [52, 83]. Besides, based on the heuristics that training with highly disparate data samples may produce better models, the Euclidean distance between a data sample of interest and centroid of all available samples at the MEC server, was considered as the data importance metric in [86]. The concept of data sample importance was also leveraged for edge robotic systems in [87], where a lightweight HARVESTNET was developed to determine whether to cache and transfer a training sample to a remote server for model re-training.
@@
In contrast to conventional wireless communication systems, energy-efficient transmission schemes for centralized edge learning should consider the data sample importance. In [88], a multi-criteria training data subset (MCTS) algorithm for prioritizing data transmission in a predictive maintenance application was developed, where the high-quality data samples are protected against wireless signal attenuation with more resources. Besides, data compression before transmission was proposed for centralized edge learning, where the compression ratio is determined according to the training loss in [84] and via an error-bounded compression scheme in [89]. To avoid the complex baseband processing at EUDs, analog modulation was exploited in [52, 83, 86] for wireless transmission, where data samples are directly mapped to channel symbols and thus vulnerable to noise corruption. As a result, a retransmission mechanism was developed in [83] to eliminate negative effects of noise, where a data sample is retransmitted until the effective signal-to-noise ratio (SNR) is above a data importance-adaptive threshold. With multiple EUDs, it is also necessary to schedule the most appropriate ones to transmit, which minimizes the communication overhead for a target learning performance [52, 86].
@@
Apart from sample-level importance, the amount of data used for model training also has a great impact on the learning performance. Although the exact relationship between data volume and training loss is hard to find, data-driven methods are able to establish power-law models for approximate characterizations [72]. By adopting a nonlinear function mapping the data volume to the training loss, a learning-centric power allocation problem for centralized edge learning was solved in [90] to decide the amount of data samples used for the training of multiple classification models. This study was extended in [91] and [92], respectively incorporating wireless resource allocation and multi-device scheduling. More realistic scenarios accounting for continuous data arrivals and limitations of MEC servers were investigated in [93, 94].
@@ III-C Centralized Edge Learning with Small Datasets
1) Data Augmentation: Data augmentation (DA) refers to the process of enhancing the size, quality, and diversity of a training dataset [95], hoping to obtain a more accurate model than using the original dataset. For instance, basic spatial operations such as rotation and noising, can be applied to images for DA. In addition to handcrafted methods, generative adversarial network (GAN) based image synthesis [96] becomes a new favorite to increase the robustness and generalizability of DA [97, 98]. DA methods for other data types are less investigated, despite with recent advances on speech recognition [99], video analytics [100], and NLP [101].
@@
DA has found its adoptions in a variety of edge AI applications. In [102], the feasibility of generating sequential IoT data via GAN for classification was validated, even with a small volume of training data. The hybrid smart medical architecture for electrocardiogram diagnostics developed in [103] adopts a similar approach to overcome the class imbalance issue in the training dataset. For WiFi signal-based human activity recognition, non-learning based data synthesis was utilized in [104] to mitigate the negative impacts caused by human movement and subject-specific bias. We are aware that while [102, 103, 104] did not directly address the energy consumption of EUDs, DA improves the learned model performance in the presence of a small and possibly low-quality training dataset. In other words, the amount of data transmitted from EUDs to the MEC server and their quality can be reduced to achieve the target learning performance with DA, thereby saving the sampling and communication energy.
2) Transfer Learning: The objective of transfer learning (TL) is to reuse the learned knowledge for a new but related model training task [105]. An attractive benefit of knowledge transfer techniques for centralized edge learning is the reduced requirement of labeled training data [106], therefore minimizing the cost of training data acquisition, including the sensing and communication energy. Besides, since MEC environments and data distributions may experience regular drifts, AI models need periodical fine-tuning [107, 108]. This can also be assisted with TL to avoid re-training models entirely from scratch on large amounts of labeled data.
The developments of industrial IoT applications have been actively utilizing TL techniques. The use of TL for industrial IoT component recognition was investigated in [109], where a pre-trained VGG-16 network on ImageNet was transferred to a new scenario with a small target domain T-Less dataset of texture-less objects. Compared with training a convolutional neural network (CNN) model from scratch, knowledge transfer reduces both the required training data volume and training time, while attaining better training accuracy. For edge camera image recognition in unmanned stores, the experimental results in [110] showed that with an elite-instance-based matching approach that appropriately selects the images to be used for TL, up to $70\%$ of source image samples can be eliminated from transmission to the target cameras. TL was also utilized for device-free crowd counting based on the WiFi channel state information (CSI) [111]. In situations where data acquisition is extremely costly and energy-expensive, some data categories may only have a few samples for centralized model training at the MEC server, for which, few-shot learning approaches [112] can be leveraged. Sample edge AI applications include object detection [113], text sentiment analysis [114], and IoT traffic classification [115].
@@ III-D Takeaways
This section summarizes energy-efficient data acquisition methods for centralized edge learning, which were developed following either of the two approaches: 1) To improve the EE of data sampling and transmission; 2) To improve the learning performance with only a small training dataset available at the MEC server. For the former approach, tractable models relating various decision variables, e.g., data sampling frequency, resolution, and context, with the learning performance are essential, which can be obtained via offline data-fitting or online learning. Although their effectiveness has been observed, existing works are mainly based on heuristics without a strong theoretical support. Besides, accurate energy models of sensing modules on EUDs are necessary to formulate valid optimization problems. For cameras, the energy model developed for visual computing systems in [74, 75] can be used. Also, a power model suitable for three-dimensional depth sensors (e.g., LiDAR cameras) is available in [116]. Moreover, in-depth studies to unveil the impacts of deploying more EUDs for sensing on EE should be conducted. The second approach stems from the fields of transfer learning and few-shot learning, while other techniques, such as meta-learning-based methods, may also be applied [105, 112]. To further reduce the communication energy consumption, it is important to understand the minimum amount of training data that needs to be transmitted from EUDs to the MEC server with the joint aid of DA and TL techniques. Since the three classes of techniques reviewed respectively in Section III-A, III-B, and III-C were examined separately, their composite effects on achieving energy-efficient data acquisition for centralized edge learning deserve further investigations.
@@ IV Energy-Efficient Cooperative Training at the Edge Network
Centralizing the model training process at an MEC server necessitates substantial communication bandwidth to transfer large volumes of training data, a requirement that is not always feasible in wireless edge networks. This approach may also infringe upon data privacy regulations [117] as local data may be private and confidential. However, advancements in embedded hardware technologies have empowered a growing population of EUDs to perform on-device model training [118, 119]. As such, cooperative model training at the edge network (a.k.a. cooperative edge learning), which distributes a model training task among multiple EUDs, has been extensively explored to leverage both the ubiquity of big data and computational resources. In this section, we focus on federated learning [120], a de facto standard for privacy-preserving cooperative training, and summarize the energy-efficient design approaches in Fig. 4. We remark that these approaches for FL are also applicable to other cooperative training paradigms at the edge network, such as decentralized edge learning (DEEL) [121] and hierarchical federated edge learning (H-FEEL) [122], which are outlined in Section IV-D.
@@ IV-A Background of Federated Learning
FL was first introduced by Google in 2016 [123]. In this framework, a set of client nodes (e.g., EUDs) collaborate to train an ML model under the coordination of a parameter server (PS), as illustrated in Fig. 5. During each training iteration, clients initially download the global model from the PS. Subsequently, each client independently updates the downloaded model using its local data, typically through methods such as stochastic gradient descent (SGD). Following this, the local model updates are uploaded to the PS for global aggregation. As the training data remains local to the clients, data privacy is significantly better protected compared to the centralized edge learning paradigm discussed in Section III.
Figure 5: A typical federated learning system (Left) and its training procedures (Right).
Although initial FL systems utilized cloud-based PSs, there has been a growing interest in deploying FL at the resource-constrained mobile edge network, a concept known as federated edge learning (FEEL). This approach holds potential for facilitating rapid cooperative training [43, 124] with an MEC server being designated as the PS. Given that FEEL includes multiple iterations of local training (e.g., backpropagation) and model uploading, the learning performance is predominantly associated with the energy consumption of both the computation and communication operations. Therefore, techniques to maximize the EE of FEEL mostly optimize either or both of the local training and model transmission processes without causing significant degradation to learning performance. In the following two sections, we introduce two categories of techniques as potential pathways to realize energy-efficient FEEL, including model and gradient compression (Section IV-B), and resource management (Section IV-C). We discuss these techniques from an EE perspective, although some of them are also essential for communication-efficient FEEL [39].